OpenStudy (anonymous):

Can someone eexplain to me how to put 148.675 to degrees minutes and seconds?

OpenStudy (wolf1728):

148.675 that's degrees right?

OpenStudy (anonymous):

Yessir.

OpenStudy (wolf1728):

148 degrees That leaves .675 multiply .675 * 60 = 40.5 minutes make that 40 minutes with ,5 left over multiply .5*60 = 30 seconds so you have 148 degrees 40 minutes 30 seconds

OpenStudy (anonymous):

Oh wow, you're the real MVP.

OpenStudy (wolf1728):

okay well thanks

OpenStudy (anonymous):

@wolf1728 What if it was the other way around? 313 31' 42''

OpenStudy (mathstudent55):

Then start from the end: 42" = 42/60 ' = 0.7' Now you have 313 deg and 31.7' Then 31.7' = 31.7/60 deg = 0.528333... deg In degrees, you have 313+0.528333... deg = 313.528333... deg